### v1

# SplatSuRe: Selective Super-Resolution for Multi-view Consistent 3D Gaussian Splatting

Abstract
3D Gaussian Splatting (3DGS) enables high-quality novel view synthesis, motivating interest in generating higher-resolution renders than those available during training. A natural strategy is to apply super-resolution (SR) to low-resolution (LR) input views, but independently enhancing each image introduces multi-view inconsistencies, leading to blurry renders. Prior methods attempt to mitigate these inconsistencies through learned neural components, temporally consistent video priors, or joint optimization on LR and SR views, but all uniformly apply SR across every image. In contrast, our key insight is that close-up LR views may contain high-frequency information for regions also captured in more distant views, and that we can use the camera pose relative to scene geometry to inform where to add SR content. Building from this insight, we propose SplatSuRe, a method that selectively applies SR content only in undersampled regions lacking high-frequency supervision, yielding sharper and more consistent results. Across Tanks & Temples, Deep Blending and Mip-NeRF 360, our approach surpasses baselines in both fidelity and perceptual quality. Notably, our gains are most significant in localized foreground regions where higher detail is desired.
Figure 1: SplatSuRe trains a 3D Gaussian Splatting [kerbl3Dgaussians] model to produce sharp, high-resolution novel views from low-resolution inputs. By selectively leveraging high-frequency cues already present in low-resolution training views and applying super-resolution only where needed, our method delivers greater detail and multi-view consistency than prior approaches without any additional training.
1 Introduction
Novel view synthesis aims to render unseen viewpoints given a set of multi-view images. 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] enables real-time, photorealistic novel view synthesis by representing scenes as an explicit set of anistropic Gaussians optimized through differentiable splatting. While 3DGS excels at efficiency and reconstruction quality, its performance is tightly coupled to the resolution of the training images. Models trained on low-resolution (LR) inputs lack access to high-frequency signals present in high-resolution (HR) views, resulting in blurry textures, over-smoothed surfaces, and aliasing artifacts when rendered at higher test-time resolutions [Yu2024MipSplatting].
When restricted to only LR views, a natural strategy is to apply super-resolution (SR) to enhance them before fitting a 3D model [feng2024srgssuperresolution3dgaussian]. However, single-image SR operates independently on each view and frequently introduces view-dependent hallucinated textures [wang2024exploiting]. When these inconsistent SR predictions are used as direct supervision, the resulting 3D optimization receives conflicting gradients across viewpoints, degrading model quality. Prior methods attempt to mitigate these inconsistencies through learned neural components, temporally consistent video priors, or joint optimization on LR and SR views [Shen2024SuperGaussian, wan2025s2gaussian, xie2024supergssuperresolution3dgaussian, huang2024assrnerfarbitraryscalesuperresolutionvoxel, yu2024gaussiansr]. However, these methods inject SR content uniformly across the image, regardless of whether a region actually benefits from generative detail or is already well-constrained by existing LR observations.
In contrast, the key observation motivating our work is that images of a scene do not sample 3D content uniformly. A low-resolution view captured up close often contains enough high-frequency detail to supervise rendering of more distant views that observe the same region only coarsely. This disparity in multi-view sampling implies that many views already receive high-frequency supervision from closer views, whereas others that do not have any closer views providing sufficient higher-frequency information would benefit from SR guidance. Applying SR indiscriminately therefore introduces unnecessary inconsistencies in well-resolved regions.
Based on this observation, we propose SplatSuRe, a selective super-resolution framework for multi-view consistent 3D Gaussian Splatting. Rather than uniformly enhancing all pixels, SplatSuRe identifies 3D regions that lack high-frequency observations and injects SR only where it is beneficial. We first compute a Gaussian fidelity score that measures how well each Gaussian is sampled across training views, then render per-view super-resolution region selection weight maps that highlight undersampled areas while suppressing SR where LR supervision is already reliable. These maps modulate SR supervision during training, allowing the model to exploit generative detail in under-resolved regions while maintaining consistency elsewhere. Through this geometry-aware selective refinement, SplatSuRe produces sharper reconstructions and improved perceptual quality without introducing additional neural components or modifying the underlying 3DGS pipeline.
In summary, we propose the following contributions:
  1. 1.
    A per-Gaussian fidelity score that quantifies how well it is resolved across views, leveraging LR geometry to estimate available high-frequency information.
  2. 2.
    A per-view spatial map of available frequency information computed using the Gaussian fidelity score that modulates SR supervision during optimization.
  3. 3.
    A selective SR training framework that jointly optimizes a 3DGS model using LR and SR supervision, injecting generative detail only where needed while preserving multi-view consistency and achieving state-of-the-art results across a diverse range of scenes.
Figure 2: Overview of our SplatSuRe framework. A high-resolution (HR) 3D Gaussian Splatting (3DGS) model is trained using low-resolution (LR) and super-resolution (SR) inputs. We first train a 3DGS model on LR inputs to identify undersampled regions and render per-view weight maps that indicate where SR is needed. During training of the HR 3DGS model, the images produced by the frozen single-image super-resolution (SISR) model are spatially weighted by these maps to form the SR loss $\mathcal{L}_{SR}$. A complementary LR loss $\mathcal{L}_{LR}$ compares the downsampled HR render against the original LR ground truth to provide consistent supervision across the entire image.
2 Related Work
3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] enables real-time novel view synthesis by representing scenes as sets of anisotropic Gaussians optimized through differentiable rasterization. While it achieves high rendering efficiency, models trained with low resolution (LR) images suffer from aliasing artifacts when rendered at higher resolutions, since Gaussians optimized at coarse scales become undersampled. Mip-Splatting [Yu2024MipSplatting] mitigates this aliasing by applying scale-adaptive 3D and 2D filtering while preserving radiance energy across resolutions, similar in spirit to Mip-NeRF [barron2022mipnerf360]. While achieving significant improvement over vanilla 3DGS, its rendering quality is still tied to the resolution of training views and blurring can occur at higher resolutions. In contrast, our method uses super-resolution to further inform high-resolution information.
Super-resolution (SR) has long been applied to neural radiance fields [mildenhall2021nerf] to enhance novel view synthesis quality [wang2022nerf, 10205402, feng2023zssrtefficientzeroshotsuperresolution, vishen2025advancingsuperresolutionneuralradiance, zheng2025supernerfganuniversal3dconsistentsuperresolution, huang2024assrnerfarbitraryscalesuperresolutionvoxel, roessle2023ganerf]. More recently, SR has been extended to 3D Gaussian Splatting (3DGS). Several variants aim to recover high-resolution detail and improve multi-view consistency [feng2024srgssuperresolution3dgaussian, Shen2024SuperGaussian, wan2025s2gaussian, xie2024supergssuperresolution3dgaussian] through residual feature learning [xie2024supergssuperresolution3dgaussian], uncertainty modeling [wan2025s2gaussian, xie2024supergssuperresolution3dgaussian], per-scene refinement [wan2025s2gaussian] or video super-resolution [Shen2024SuperGaussian]. Among these, SRGS [feng2024srgssuperresolution3dgaussian] jointly optimizes Gaussian parameters using both LR ground-truth images and super-resolved views produced by a frozen single-image SR model. However, applying this enhancement uniformly across all regions does not eliminate the effect of inconsistencies introduced by super-resolution. S2Gaussian [wan2025s2gaussian] focuses on sparse view reconstruction and proposes an inconsistency modeling module trained per-scene to reduce inconsistencies in SR images. SuperGaussian [Shen2024SuperGaussian] applies a video SR network to frames rendered from a low-resolution 3DGS model, using the resulting temporally consistent sequence for retraining the 3D model. While these approaches improve fidelity, they either rely on additional neural components to enforce consistency or lack spatial adaptivity in the use of SR. In contrast, our method utilizes camera pose information relative to scene geometry to determine undersampled regions, and selectively applies super-resolution in those regions to enhance sharpness while maintaining fidelity.
Diffusion methods offer complementary advances in 3D representation and enhancement tasks. Several works jointly optimize diffusion and 3D parameters for view-consistent generation and super-resolution [lin2025diffsplat, DiffGS, yu2024gaussiansr, chen2025bridgingdiffusionmodels3d]. GaussianSR [yu2024gaussiansr] distills information from 2D super-resolution models as a loss for training a 3DGS model. 3DSR [chen2025bridgingdiffusionmodels3d] uses a 3DGS model to enforce 3D consistency in the diffusion process, fitting it multiple times through denoising diffusion steps. Diffusion priors have also been applied as post-processing to enhance renders from 3DGS models [liu20243dgs, wu2025difix3d, gsfix3d]. While these methods rely on large pretrained diffusion models to predict images that reduce 3D inconsistencies, we leverage explicit geometric relationships between cameras and scene structure to determine where generative detail is needed, which can then be added via any of these complementary methods.
In addition to image and video-based models, SR can also be performed directly in 3D, circumventing multi-view inconsistencies. Geometric point-based networks [yu2018pu, Liu_2024_CVPR, fang2024egp3dedgeguidedgeometricpreserving, dinesh20193dpointcloudsuperresolution] locally upscale point clouds to recover fine-grained geometry. However, these approaches typically only upsample geometry, while texture resolution still relies on 2D methods. As before, our method is orthogonal to these techniques and can be combined with them to further improve reconstruction fidelity.
3 Background
3.1 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] models a scene as a set of 3D Gaussians, each represented by a mean position $\boldsymbol{\mu}\in\mathcal{R}^{3}$, per-axis scale $\boldsymbol{s}\in\mathcal{R}^{3}$, rotation vector $\boldsymbol{q}\in\mathcal{R}^{4}$, scalar opacity $o\in\mathcal{R}_{+}$, and view-dependent color $\boldsymbol{c}$, represented as a base color with spherical harmonic coefficients. Volumetric rendering with alpha blending is used to splat the Gaussians onto the image plane and the resulting color for pixel $\boldsymbol{p}$ is given by:
$C(\boldsymbol{p})=\sum_{i=1}^{N}\boldsymbol{c}^{i}\alpha^{i}(\boldsymbol{p})\prod_{j=1}^{i-1}(1-\alpha^{j}(\boldsymbol{p})),$ (1)
where $N$ is the number of Gaussians intersecting the pixel ray, $\alpha^{i}(\boldsymbol{p})=o^{i}e^{-\frac{1}{2}(\boldsymbol{p}-\boldsymbol{\mu}^{i})^{T}(\boldsymbol{\Sigma}_{{2D}}^{i})^{-1}(\boldsymbol{p}-\boldsymbol{\mu}^{i})}$ is the contribution of the $i^{th}$ Gaussian, and $\boldsymbol{c}^{i}$ is its view-dependent color. The 2D covariance matrix $\boldsymbol{\Sigma}_{2D}$ is given by the EWA Splatting approximation [zwicker2002ewa]. Gaussians are sorted in depth order before splatting to ensure that transmittance is computed correctly.
The model is initialized using Structure from Motion (SfM) on the training views to provide camera parameters and produce a sparse point cloud, serving as the initial Gaussian means. During model training, images are randomly selected from the training set, and a weighted sum of $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{D-SSIM}}$ losses is used to optimize the Gaussian parameters using gradient descent. Gaussians are split and cloned throughout training to increase scene fidelity and ensure sufficient primitives where needed. We retain the standard 3DGS rendering and densification pipeline, modifying only the supervision losses for our method.
3.2 Super-Resolution for Gaussian Splatting
When trained solely on low-resolution (LR) images, 3DGS models lack access to the high-frequency cues present in high-resolution (HR) views, leading to over-smoothed textures and incomplete fine structure. This limitation has motivated the integration of super-resolution (SR) models into the 3DGS training pipeline. SRGS [feng2024srgssuperresolution3dgaussian] employs a frozen single-image SR model to generate super-resolved views and jointly optimizes Gaussian parameters against both LR ground truth and SR outputs, supplying explicit high-frequency supervision that LR views alone cannot provide. However, SRGS applies SR uniformly across the entire image, even in regions that already receive reliable high-frequency supervision from nearby LR views. Since SR is applied independently to each image, the generated details are not necessarily multi-view consistent – injecting them everywhere can introduce geometric or texture inconsistencies, leading to averaging effects in the model that render as blurring.
These observations reveal a fundamental challenge: super-resolution is not uniformly beneficial across the scene. As illustrated in Figure 3, some regions already receive sufficient high-frequency supervision from closer LR training views, so adding generated detail introduces unnecessary inconsistencies that harm cross-view coherence. Other regions, particularly those that are distant or sparsely observed, are undersampled and require SR to recover missing details. This motivates a selective, geometry-aware strategy that determines where SR should influence optimization. Instead of treating all pixels equally, we aim to exploit the multi-view sampling pattern of each Gaussian to determine which regions are sufficiently constrained by LR observations and which require additional SR guidance.
4 Method
Super-resolution (SR) mostly benefits regions that lack reliable high-frequency supervision, motivating a geometry-aware mechanism for deciding where SR should influence 3DGS optimization. Our method identifies these undersampled areas and applies SR only where high-frequency detail is missing, improving sharpness while avoiding unnecessary inconsistencies. To achieve this, we compute a Gaussian fidelity score that measures how well each Gaussian is sampled across training views, then render per-view weight maps that highlight undersampled areas while suppressing SR where LR supervision is already reliable. Incorporating these weight maps into a combined LR–SR objective yields our SplatSuRe framework, shown in Figure 2, which selectively injects detail where it is beneficial while preserving multi-view consistency elsewhere.
4.1 Gaussian Fidelity Score
Images capturing a scene do not contribute equal amounts of high-frequency information for 3D reconstruction. A low-resolution (LR) view taken at a short distance or with a long focal length can contain more fine detail than a high-resolution (HR) view captured from further away. We leverage this inherent disparity to determine which 3D regions already have adequate high-frequency supervision and which require additional details from super-resolution (SR). Figure 3 illustrates how a nearby LR image can provide the high-frequency detail needed to supervise a distant viewpoint.
To measure per-Gaussian relative sampling frequency, we first train a low-resolution 3DGS model using the LR images. This provides stable scene geometry and allows us to compute each Gaussian’s screen-space radius in every training view. Following 3DGS, the screen-space radius, measured in pixel units, of Gaussian $\mathcal{G}^{i}$ is:
$r^{i}=3\sqrt{\text{max}(\lambda^{i}_{1},\lambda^{i}_{2})},$ (2)
where $\lambda^{i}_{1}$ and $\lambda^{i}_{2}$ are the eigenvalues of the Gaussian’s 2D covariance matrix $\boldsymbol{\Sigma}^{i}_{2D}$, calculated as:
$\lambda^{i}_{1},\lambda^{i}_{2}=$ $\frac{1}{2}\text{tr}(\boldsymbol{\Sigma}^{i}_{2D})\pm\sqrt{\text{max}\{0.1,\frac{1}{4}\text{tr}^{2}(\boldsymbol{\Sigma}^{i}_{2D})-|\boldsymbol{\Sigma}^{i}_{2D}|\}},$ (3)
[…]
$r^{i}_{min}=\min_{t\in T}{r^{i}_{t}},\qquad r^{i}_{max}=\max_{t\in T}{r^{i}_{t}},$ (4)
$\rho^{i}=r^{i}_{max}/r^{i}_{min}.$ (5)
A high ratio indicates that the Gaussian is sampled at varied frequencies across views, meaning some views observe it with high fidelity and can supervise the others. Conversely, a ratio close to one indicates uniformly sampling across views, indicating that this region lacks any higher-frequency observations and requires SR to add generated details. This interpretation also holds for view regions where the Gaussian projects near its maximal radius, since no other views provide higher frequency information.
Figure 3: Disparity in high-frequency ground truth information across different views. Low-resolution ground truth from near cameras provides high-resolution information for rendering distant views, reducing the need for additional generated detail in those views. Conversely, super-resolution is needed in views where no other camera provides higher-resolution information.
Note that 3DGS dilates each Gaussian by convolving it with a fixed low-pass Gaussian filter to prevent aliasing and ensure a minimal rendering size:
$\boldsymbol{\Sigma}^{i}_{2D}=\boldsymbol{\Sigma}^{i}_{2D}+s\boldsymbol{I},$ (6)
where s=0.3s{=}0.3 is the amount of dilation. Since this dilation artificially inflates the radius, especially for distant Gaussians where the blur dominates, we exclude this dilation when computing radii for the ratio.
We then transform the raw ratio $\rho^{i}$ into a Gaussian fidelity score that maps each Gaussian to a weight in $[0,1]$, with lower values indicating a greater need for SR. The ratio is first offset by a threshold $\tau$ and then mapped into the unit interval using a sigmoid function, giving the per-Gaussian score:
$\text{score}_{\mathcal{G}^{i}}=\sigma(\frac{\rho^{i}-\tau}{k}),$ (7)
where $\sigma$ is the sigmoid function and k=0.05k{=}0.05 controls the smoothness of the transition from 0 to 1. The ratio threshold $\tau$ is a hyper-parameter that depends on the structure of the scene and the consistency of the SR model across views. Scores for Gaussians visible in fewer than three views are set to zero because these regions are poorly constrained. Higher scores correspond to Gaussians that are already well-captured by LR supervision, while lower scores identify regions where SR should be applied more heavily.
4.2 Super-Resolution Region Selection
Our Gaussian fidelity score provides a scene-level measure of how well a Gaussian is observed across views, but supervision of high-resolution model updates requires pixel-wise weight maps for each training view. For a given training view $t$, we identify the set of Gaussians whose maximal radius occurs in its rendered view:
$\mathcal{M}(t)$ $=\{\mathcal{G}^{i}\ |\ t=\operatorname*{argmax}_{t\in T}r^{i}_{t}\;\forall i\in N\},$ (8)
where $\mathcal{G}^{i}$ is the $i^{th}$ Gaussian, $r^{i}_{t}$ is its screen-space radius, and $N$ is the total number of Gaussians. These are Gaussians that do not receive higher-frequency information from another view. The weight map for training view $t$ is then rendered as:
$W^{\prime}_{t}$ $=(1-\text{Render}(\mathbf{score}_{\mathcal{G}}))+\text{Render}(\mathbf{1}_{\mathcal{M}(t)}(\mathcal{G})),$ (9)
where Render()($\cdot$) denotes splatting the LR model by alpha-blending the specified per-Gaussian values, $\mathcal{G}$ is the set of all Gaussians, $\mathbf{score}_{\mathcal{G}}$ is the vector of Gaussian fidelity scores, and $\mathbf{1}_{\mathcal{M}(t)}(\mathcal{G})$ is an indicator that is 1 for Gaussians in $\mathcal{M}(t)$ and 0 otherwise. The first term in Equation 9 ensures that undersampled regions with low fidelity scores receive SR, while the second term ensures that areas observed most closely by the current view also receive SR because no alternative view provides higher resolution information. Finally, the weight map is normalized to ensure the magnitude of the SR loss is consistent across views. As illustrated in Figure 4, this per-view weight map is high in regions requiring SR and low in regions already sufficiently supervised by LR views.
Figure 4: Super-resolution weight maps. Bright regions indicate areas where generative detail is required, while dark regions correspond to areas well-sampled by other low-resolution views. Note that high weights are obtained in regions that are either not sampled closely, such as background trees behind the tractor, or where other views do not provide higher resolution information, such as the foreground table in the ballroom.
4.3 SplatSuRe Training Objective
Our super-resolution region selection method determines how much SR should influence each pixel during optimization. Rather than uniformly applying SR to the entire image during training, we incorporate our weight map $W_{t}$ to supervise training using two complementary signals: (1) low-resolution ground truth images, and (2) selectively-weighted super-resolved images.
In each training iteration, the model is rendered at the target high resolution. The render $R_{HR}$ is downsampled to produce $R_{HR}\downarrow$ and compared with the LR ground truth $I_{LR}$, yielding the LR loss:
$\mathcal{L}_{LR}=(1-\lambda)\mathcal{L}_{1}(R_{HR}\downarrow,I_{LR})+\lambda\mathcal{L}_{\text{D-SSIM}}(R_{HR}\downarrow,I_{LR}),$ (10)
where $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{D-SSIM}}$ are the losses used in 3DGS. The super-resolved image $I_{SR}$ is compared with the HR render $R_{HR}$ using a spatially weighted loss:
$\mathcal{L}_{SR}=(1-\lambda)\mathcal{L}_{1}^{W}(R_{HR},I_{SR})+\lambda\mathcal{L}_{\text{D-SSIM}}^{W}(R_{HR},I_{SR}),$ (11)
where each pixel’s contribution is scaled by its weight in $W_{t}$. High weights amplify SR supervision in undersampled regions, while low weights suppress it where LR views already provide reliable high-frequency information.
The overall objective combines LR ground truth supervision with selectively-weighted SR guidance:
$\mathcal{L}=(1-\gamma)\mathcal{L}_{LR}+\gamma\mathcal{L}_{SR},$ (12)
where $\gamma$ controls the relative contribution of each term. This SplatSuRe formulation produces a high-resolution 3DGS model that leverages SR only where it improves reconstruction quality, avoiding inconsistencies in regions already well-constrained by LR views.
Figure 5: Qualitative results on Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018] and Mip-Nerf 360 [barron2022mipnerf360]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [Yu2024MipSplatting] and SRGS [feng2024srgssuperresolution3dgaussian], our method produces sharper, more faithful reconstructions that better align with ground truth while maintaining cross-view consistency. It preserves finer details in text (red box on truck), high-frequency patterns (yellow box on carpet and green box on tray) and recognizable distant objects observed in other views (blue box on church mural). Notably, it reduces Gaussian artifacts (orange arrow) observed in other methods. Additional results in Appendix A.3.
5 Experiments
We integrate our SplatSuRe method into the 3D Gaussian Splatting (3DGS) codebase [liu20243dgs], adding modules for true Gaussian radius computation, weight map computation and rendering, and auxiliary loss terms. Experiments are conducted at $4\times$ super-resolution using StableSR [wang2024exploiting] and using a ratio threshold $\tau{=}1.1$ to generate weight maps. Our choice of ratio threshold and SR model are ablated in Sections 7.1 and 7.2, respectively. We use $\lambda{=}0.2$ and $\gamma{=}0.4$ for our losses defined in Equations 10-12.
Datasets.
[…]
Table 1: Quantitative results on Tanks & Temples [Knapitsch2017tandt]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. The best, second best and third best entries are highlighted. Our SplatSuRe method achieves the strongest results across all metrics.
Method Tanks & Temples [Knapitsch2017tandt]
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ CMMD $\downarrow$ DreamSim $\downarrow$ MUSIQ $\uparrow$ NIQE $\downarrow$
3DGS (LR) [kerbl3Dgaussians] 0.669 19.41 0.350 71.58 2.013 0.0895 57.776 3.412
StableSR [wang2024exploiting] 0.751 22.47 0.300 59.29 1.123 0.0667 56.748 4.945
Mip-Splatting [Yu2024MipSplatting] 0.767 23.10 0.303 52.46 1.137 0.0597 46.571 5.043
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.771 23.32 0.286 49.11 1.048 0.0535 55.209 4.633
Ours + StableSR [wang2024exploiting] 0.784 23.81 0.272 37.72 1.040 0.0413 58.332 3.928
Table 2: Quantitative results on Deep Blending [DeepBlending2018] and Mip-NeRF 360 [barron2022mipnerf360]. Our SplatSuRe method achieves the strongest results across all metrics on Deep Blending and outperforms SRGS [feng2024srgssuperresolution3dgaussian] on Mip-NeRF 360. Appendix A.1 and A.2 present $8\times$ SR and per-scene results.
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ CMMD $\downarrow$ DreamSim $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ CMMD $\downarrow$ DreamSim $\downarrow$
3DGS (LR) [kerbl3Dgaussians] 0.836 26.72 0.335 0.868 0.0574 0.635 20.67 0.384 0.645 0.0529
StableSR [wang2024exploiting] 0.845 27.16 0.325 0.630 0.0505 0.699 24.28 0.351 0.725 0.0387
Mip-Splatting [Yu2024MipSplatting] 0.865 28.43 0.327 0.690 0.0398 0.759 26.48 0.292 0.183 0.0132
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.861 28.23 0.317 0.630 0.0409 0.734 25.92 0.317 0.339 0.0194
Ours + StableSR [wang2024exploiting] 0.872 29.01 0.306 0.496 0.0330 0.740 26.34 0.323 0.339 0.0179
Baselines.
We compare SplatSuRe with four representative baselines using their official implementations on our evaluation datasets. 3DGS (LR) [kerbl3Dgaussians] is trained on low-resolution inputs and serves as the primary baseline. StableSR [wang2024exploiting] applies a single-image SR model to the low-resolution training data and fits a 3DGS model on the super-resolved images. Mip-Splatting [Yu2024MipSplatting] is trained at low resolution and mitigates aliasing through 3D and 2D multi-scale filtering for high-resolution rendering. Finally, SRGS [feng2024srgssuperresolution3dgaussian] optimizes Gaussian parameters using both low-resolution and super-resolved images from a frozen SR model, without additional pretraining or scene-specific fine-tuning.
Metrics.
We evaluate both reconstruction fidelity and perceptual realism using eight complementary metrics. For reference-based assessment, we report SSIM, PSNR, and LPIPS [zhang2018perceptual], which measure pixel and feature-level agreement with ground-truth high-resolution images. To capture distributional and semantic perceptual quality, we include FID [heusel2017gans], CMMD [jayasumana2024rethinkingfidbetterevaluation], and DreamSim [fu2023dreamsim], which compare the feature distributions or embeddings of generated and real images in pretrained perceptual spaces. Finally, we evaluate no-reference perceptual quality using MUSIQ [ke2021musiqmultiscaleimagequality] and NIQE [6353522], which estimate realism and naturalness without ground-truth supervision. While PSNR and SSIM directly measure pixel fidelity at the native resolution, the perceptual and distributional metrics internally downsample or resize images before feature extraction, reducing sensitivity to high-frequency details. As a result, improvements in fine-scale sharpness – central to super-resolution methods – may be underrepresented by these metrics, motivating our use of a broad set of complementary evaluations.
6 Results
Qualitative Results.
Figures 1 and 5 present representative renderings from our method. Our approach produces consistently sharper reconstructions while preserving smoothness in uniformly textured regions. In contrast, Mip-Splatting yields overly blurred results because it is trained only on LR and its 3D anti-aliasing filter excessively blurs all Gaussians. SRGS recovers higher-frequency content in some areas but remains constrained by the underlying SR images – regions where the SR output is consistent appear sharp, whereas areas with view-inconsistent SR predictions become noticeably blurred or distorted.
Quantitative Results.
Table 1 presents results on Tanks & Temples [Knapitsch2017tandt]. Our method achieves the highest image quality across nearly all metrics, demonstrating both its effectiveness and robustness. The only exception is NIQE, where 3DGS attains a better score. Similarly, SRGS ranks second across most reference and perceptual metrics but lags behind in MUSIQ and NIQE, where 3DGS performs better. This trend arises because the aliased, noisy renderings of low-resolution 3DGS coincidentally resemble the natural image statistics favored by no-reference quality metrics, whereas the smoother and more coherent outputs of Mip-Splatting and SR-based methods appear less “natural” under such measures. In contrast, our approach delivers substantially higher perceptual fidelity and cross-view consistency, producing sharper and more realistic high-resolution reconstructions that align closely with both human perception and ground-truth images.
Figure 6: Effect of ratio threshold on Tanks & Temples [Knapitsch2017tandt]. Weight maps, where bright regions indicate higher SR influence, are shown below the corresponding ratio thresholds. $\tau{=}0$ and $\tau{=}\infty$ correspond to zero and full use of super-resolution. SR is initially helpful in improving rendering quality, but excessive use worsens results. The effect of ratio threshold on different scenes is analyzed in Appendix A.2.
Table 2 reports results on Deep Blending [DeepBlending2018] and Mip-NeRF 360 [barron2022mipnerf360]. Our method achieves the strongest performance across all metrics on Deep Blending. On Mip-NeRF 360, it outperforms SRGS on every metric except LPIPS; however, both SR-based approaches are surpassed by Mip-Splatting. This advantage arises from the characteristics of Mip-NeRF 360, where smooth, circular camera trajectories provide dense multi-view coverage and minimal undersampling. In such well-sampled settings, Mip-Splatting’s multi-scale anti-aliasing filters effectively preserve radiance energy and suppress aliasing, whereas SR-based methods may introduce slight inconsistencies when enhancing already well-resolved regions. Additionally, the LR images in Mip-Nerf 360 are approximately twice as large as those in the other datasets. Since most high-frequency details are already preserved at this scale, SR offers minimal improvement. Overall, SplatSuRe achieves the highest performance on Tanks & Temples and Deep Blending and outperforms SRGS on Mip-NeRF 360. Results for $8\times$ SR, additional visualizations, and per-scene metrics for Tables 1 and 2 can be found in Appendix A.1, A.3, and A.4.
7 Ablations
7.1 Ablation on Ratio Threshold
Scenes with different geometries and scales exhibit varying distributions of high and low-frequency content across training views. To account for this variability, we examine the effect of the ratio threshold used to determine where SR is applied. As shown in Figure 6, introducing a small amount of SR initially improves both PSNR and LPIPS by recovering fine details in undersampled regions. However, applying SR too aggressively leads to inconsistencies across views, degrading overall performance. This trend is most pronounced in scenes with varied camera-object distances, where sampling density differs significantly across views. Based on this analysis, we select ratio threshold $\tau{=}1.1$, which provides the best trade-off between sharpness and consistency across scene, in our main experiments. See Appendix A.2 for a detailed analysis on different scenes.
7.2 Ablation on Super-Resolution Model
We evaluate our method using two single-image super-resolution (SISR) models: SwinIR[liang2021swinir] and StableSR[wang2024exploiting]. SwinIR is optimized for high fidelity under pixel-based metrics such as PSNR, producing accurate yet often over-smoothed reconstructions. In contrast, StableSR incorporates diffusion-based generative priors to synthesize sharper, more detailed, and perceptually realistic images, sometimes at the expense of lower PSNR. As shown in Table 3, our approach improves performance across both models, demonstrating that it is agnostic to the choice of SISR backbone. Notably, our gains are larger with StableSR, as its higher perceptual quality comes at the cost of multi-view inconsistencies. Across both methods, StableSR produces better perceptual quality metrics, motivating our use of StableSR for the main experiments where perceptual realism is prioritized alongside reconstruction fidelity.
Table 3: Comparison of SwinIR [liang2021swinir] and StableSR [wang2024exploiting] on Tanks & Temples [Knapitsch2017tandt]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. Our method outperforms SRGS with either model. While SwinIR achieves higher PSNR due to its conservative reconstruction, we choose StableSR for our main experiments for its superior perceptual quality.
Method SwinIR [liang2021swinir] StableSR [wang2024exploiting]
PSNR $\uparrow$ CMMD $\downarrow$ PSNR $\uparrow$ CMMD $\downarrow$
SRGS [feng2024srgssuperresolution3dgaussian] 23.84 1.137 23.32 1.048
Ours 24.06 1.135 23.81 1.043
8 Limitations and Future Work
Our method reduces multi-view inconsistencies by suppressing super-resolution in regions already well-captured by low-resolution views. However, this conservative strategy may miss cases where SR could still offer useful refinements. For example, many SR models generate stable, visually sharp detail along high-contrast boundaries. Incorporating adaptive mechanisms that selectively apply SR in such regions could further enhance reconstruction quality. Additionally, our framework currently operates at a single upsampling level, modulating only the strength of SR supervision. Extending this to a multi-scale formulation – where smaller upsampling ratios offer more consistent detail than larger ones – may offer finer control over SR integration and lead to additional gains in sharpness.
9 Conclusion
We introduced SplatSuRe, a selective super-resolution framework for 3D Gaussian Splatting that applies SR only where high-frequency information is missing. By leveraging scene geometry and camera pose to estimate per-Gaussian sampling fidelity, SplatSuRe identifies undersampled regions and modulates SR supervision through per-view weight maps that highlight where generative detail is truly needed. This geometry-aware strategy injects enhanced detail and sharpness where required while preserving multi-view consistency, achieving state-of-the-art results across a diverse range of scenes.
Acknowledgments
This work was made possible by NSF Grants 21-37229 and 22-35050, DARPA TIAMAT, and the NSF TRAILS Institute (2229885). This research is based upon work supported by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via IARPA R&D Contract No. 140D0423C0076. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of the ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Additional support was provided by Coefficient Giving.
Appendix A Appendix
A.1 Results for 8× Super-Resolution
Table 4 reports quantitative results at $8\times$ super-resolution for Tanks & Temples [Knapitsch2017tandt]. We downsample the original $1920\times 1080$ images by $16\times$ and upsample by $8\times$ to half the native resolution. Our method achieves the strongest performance across most metrics (SSIM, PSNR, LPIPS, FID and DreamSim). For MUSIQ and NIQE, 3DGS (LR) attains higher scores, consistent with the NIQE behavior discussed in Section 6. Our CMMD score ranks fourth among the compared methods. Prior work [jayasumana2024rethinkingfidbetterevaluation] shows that CMMD is sensitive to high-frequency distortions introduced by noise in the latent embedding space. We hypothesize that the sharpening effects like mild aliasing or overly crisp edges introduced at large upsampling factors may be interpreted by CMMD as distortions, even though they improve perceptual quality and are reflected positively by other metrics. Since no single metric fully characterizes visual fidelity, we report a broad suite of metrics to provide a more comprehensive assessment of our method.
[…]
A.2 Additional Ratio Threshold Analysis
Different scenes exhibit distinct behaviors as the ratio threshold and the corresponding amount of super-resolution (SR) information are increased. Most scenes benefit from a moderate amount of SR, but experience a sharp drop in image quality when it is excessively applied. Others plateau in image quality or continue improving, although with diminishing returns. We visualize these trends by plotting PSNR and LPIPS across ratio thresholds for three representative scenes from each category.
Figure 7 illustrates scenes that benefit from an optimal amount of SR. These scenes exhibit a characteristic pattern in which image quality initially improves with moderate SR, but decreases sharply when it is excessively applied. Our method identifies the most poorly sampled regions and selectively applies SR to them, yielding a substantial initial quality boost. However, applying excessive SR introduces multi-view inconsistencies that rapidly degrades image quality. This behavior appears consistently across most scenes, supporting our hypothesis that selectively applying SR is more beneficial than applying it uniformly.
Figure 8 presents scenes that plateau in image quality or continue to benefit slightly as the amount of SR is increased. Our method identifies the most poorly sampled regions and selectively applies SR to them, producing the sharp early quality gains at lower ratio thresholds. Beyond this point, increasing the amount of SR yields diminishing returns or no improvement. This is particularly observed in scenes where the input images already contain substantial high-frequency detail. In such cases, SR produces simpler sharpening or edge-enhancement effects rather than hallucinating new structure, making uniform application less harmful and sometimes marginally beneficial. This behavior is especially common in outdoor scenes in Mip-NeRF 360 [barron2022mipnerf360], where the downsampled $\sim 500\times 375$ images remain relatively high-resolution and therefore do not exhibit the multi-view inconsistencies that typically arise when excessive SR is applied.
A.3 Additional Visualizations
Figures 9 and 10 present additional qualitative results on Tanks & Temples [Knapitsch2017tandt]. Consistent with the examples in Figures 1 and 5 discussed in Section 6, SplatSuRe produces sharper reconstructions than competing methods while reducing artifacts and preserving smoothness in uniformly textured regions.
A.4 Per-Scene Metrics
Tables 6, 7, 8, 9, 10, 11, 12, and 13 report per-scene SSIM, PSNR, LPIPS, FID, CMMD, DreamSim, MUSIQ, and NIQE results on Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360] for all methods evaluated in Section 5. Across individual scenes, we observe the same trends as in the averaged results reported in Tables 1 and 2 in Section 6: SplatSuRe achieves the strongest performance on Tanks & Temples and Deep Blending, and outperforms SRGS on Mip-NeRF 360.
[…]
Method Tanks & Temples [Knapitsch2017tandt]
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ CMMD $\downarrow$ DreamSim $\downarrow$ MUSIQ $\uparrow$ NIQE $\downarrow$
3DGS (LR) [kerbl3Dgaussians] 0.537 16.40 0.473 172.44 3.795 0.2427 46.870 3.976
StableSR [wang2024exploiting] 0.664 21.49 0.406 99.80 1.878 0.1100 38.929 6.293
Mip-Splatting [Yu2024MipSplatting] 0.661 21.53 0.428 109.01 1.941 0.1183 32.417 6.645
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.674 21.97 0.399 92.91 1.891 0.0981 38.322 6.138
Ours + StableSR [wang2024exploiting] 0.692 22.48 0.378 77.17 2.007 0.0774 44.428 5.341
[…]
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ CMMD $\downarrow$ DreamSim $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ CMMD $\downarrow$ DreamSim $\downarrow$
3DGS (LR) [kerbl3Dgaussians] 0.783 24.50 0.404 2.051 0.1280 0.509 18.17 0.491 2.395 0.1337
StableSR [wang2024exploiting] 0.821 26.68 0.383 1.159 0.0705 0.630 24.23 0.445 1.096 0.0546
Mip-Splatting [Yu2024MipSplatting] 0.831 27.24 0.391 1.127 0.0689 0.656 24.80 0.420 0.610 0.0243
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.830 27.35 0.377 1.079 0.0622 0.648 24.88 0.422 0.665 0.0320
Ours + StableSR [wang2024exploiting] 0.843 28.03 0.357 0.879 0.0489 0.655 25.08 0.406 0.546 0.0231
(a) ballroom – Tanks & Temples [Knapitsch2017tandt]
(b) kitchen – Mip-NeRF 360 [barron2022mipnerf360]
(c) francis – Tanks & Temples [Knapitsch2017tandt]
Figure 7: Representative scenes that benefit from an optimal amount of super-resolution. Top: Image quality vs. ratio threshold plots. Bottom: ground truth images illustrating scene structure for (a) ballroom from Tanks & Temples [Knapitsch2017tandt], (b) kitchen from Mip-NeRF 360 [barron2022mipnerf360] and (c) francis from Tanks & Temples. Applying SR to the most poorly sampled regions yields large gains in image quality, whereas excessive SR introduces multi-view inconsistencies that sharply degrade quality. Most scenes exhibit this behavior, supporting our hypothesis that selectively applying SR is more beneficial than applying it uniformly.
(a) bicycle – Mip-NeRF 360 [barron2022mipnerf360]
(b) garden – Mip-NeRF 360 [barron2022mipnerf360]
(c) stump – Mip-NeRF 360 [barron2022mipnerf360]
Figure 8: Representative scenes that plateau in image quality or continue to benefit from increased amounts of super-resolution. Top: Image quality vs. ratio threshold plots. Bottom: ground truth images illustrating scene structure for (a) bicycle, (b) garden, and (c) stump from Mip-NeRF 360 [barron2022mipnerf360]. Applying SR to the most poorly sampled regions yields large gains in image quality, while further increasing SR provides diminishing returns or no improvement. This is particularly observed in scenes where the input images already contain substantial high-frequency detail. In such cases, SR produces simpler sharpening or edge-enhancement effects rather than hallucinating new structure, making uniform application less harmful and sometimes marginally beneficial.
Figure 9: Additional qualitative results on Tanks & Temples [Knapitsch2017tandt]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [Yu2024MipSplatting] and SRGS [feng2024srgssuperresolution3dgaussian], our method produces sharper, more faithful reconstructions that better align with ground truth while maintaining cross-view consistency. It preserves high-frequency patterns (orange boxes on statue), fine details in text (purple box on vehicle), and reduces artifacts while preserving sharpness (green boxes on curtains, blue boxes in playground).
Figure 10: Additional qualitative results on Tanks & Temples [Knapitsch2017tandt]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [Yu2024MipSplatting] and SRGS [feng2024srgssuperresolution3dgaussian], our method produces sharper, more faithful reconstructions that better align with ground truth while maintaining cross-view consistency. It preserves high-frequency geometry (pink boxes on barn) and retains sharp texture (yellow boxes on lighthouse, red boxes on horse statue, blue boxes on tank).
Table 6: SSIM $\uparrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. The best, second best and third best entries are highlighted.
Method Tanks & Temples [Knapitsch2017tandt]
auditorium ignatius palace ballroom panther barn lighthouse playground courtroom m60 temple caterpillar family train francis truck church horse museum
3DGS (LR) [kerbl3Dgaussians] 0.807 0.527 0.666 0.578 0.752 0.729 0.742 0.628 0.647 0.735 0.715 0.602 0.566 0.648 0.779 0.649 0.682 0.675 0.579
StableSR [wang2024exploiting] 0.849 0.646 0.709 0.702 0.819 0.771 0.777 0.720 0.725 0.810 0.761 0.686 0.753 0.705 0.845 0.759 0.744 0.811 0.683
Mip-Splatting [Yu2024MipSplatting] 0.838 0.681 0.687 0.729 0.847 0.790 0.789 0.746 0.730 0.836 0.762 0.704 0.788 0.733 0.855 0.786 0.753 0.827 0.689
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.861 0.671 0.713 0.734 0.846 0.794 0.789 0.742 0.746 0.835 0.773 0.703 0.780 0.729 0.857 0.784 0.761 0.829 0.704
Ours + StableSR [wang2024exploiting] 0.859 0.694 0.717 0.746 0.859 0.808 0.807 0.764 0.752 0.846 0.786 0.714 0.795 0.743 0.866 0.798 0.781 0.835 0.716
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
drjohnson playroom bicycle bonsai counter flowers garden kitchen room stump treehill
3DGS (LR) [kerbl3Dgaussians] 0.828 0.843 0.527 0.776 0.772 0.429 0.559 0.719 0.836 0.581 0.516
StableSR [wang2024exploiting] 0.837 0.854 0.604 0.848 0.833 0.497 0.654 0.751 0.858 0.669 0.571
Mip-Splatting [Yu2024MipSplatting] 0.855 0.874 0.676 0.917 0.872 0.544 0.719 0.870 0.897 0.740 0.596
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.851 0.870 0.653 0.896 0.863 0.533 0.692 0.825 0.888 0.671 0.590
Ours + StableSR [wang2024exploiting] 0.867 0.877 0.646 0.906 0.867 0.503 0.690 0.869 0.896 0.706 0.582
Table 7: PSNR $\uparrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
auditorium ignatius palace ballroom panther barn lighthouse playground courtroom m60 temple caterpillar family train francis truck church horse museum
3DGS (LR) [kerbl3Dgaussians] 22.14 17.13 18.35 18.23 22.49 22.97 20.47 19.85 19.53 21.72 18.99 19.36 15.09 18.80 22.47 17.38 20.00 16.30 17.44
StableSR [wang2024exploiting] 23.62 20.52 19.68 22.20 25.23 25.18 21.65 23.24 21.75 24.72 20.24 22.44 22.54 20.55 25.97 22.86 22.02 22.72 19.76
Mip-Splatting [Yu2024MipSplatting] 23.46 21.22 18.24 23.15 27.11 26.85 21.29 24.13 22.26 26.16 20.31 22.99 24.09 21.33 26.62 23.94 21.99 23.48 20.28
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 24.23 21.26 19.69 23.44 26.98 26.83 21.74 24.12 22.74 26.17 20.49 22.98 23.82 21.28 26.55 23.87 22.54 23.72 20.63
Ours + StableSR [wang2024exploiting] 24.66 21.61 20.24 23.82 27.38 27.33 22.82 25.12 22.99 26.63 21.39 23.34 24.00 21.79 27.34 24.11 23.18 23.94 20.77
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
drjohnson playroom bicycle bonsai counter flowers garden kitchen room stump treehill
3DGS (LR) [kerbl3Dgaussians] 26.29 27.15 18.50 22.75 23.27 17.60 18.64 19.55 25.93 20.34 19.44
StableSR [wang2024exploiting] 26.60 27.71 22.87 26.48 26.44 19.92 23.80 24.80 28.32 24.24 21.64
Mip-Splatting [Yu2024MipSplatting] 27.54 29.33 24.28 30.65 28.28 21.05 25.63 29.48 30.69 26.20 22.08
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 27.38 29.08 24.05 29.81 28.11 20.91 25.20 28.26 30.42 24.58 21.97
Ours + StableSR [wang2024exploiting] 28.46 29.57 24.04 30.54 28.39 20.55 25.31 29.56 31.05 25.61 22.02
Table 8: LPIPS $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
auditorium ignatius palace ballroom panther barn lighthouse playground courtroom m60 temple caterpillar family train francis truck church horse museum
3DGS (LR) [kerbl3Dgaussians] 0.301 0.411 0.380 0.337 0.309 0.317 0.303 0.371 0.376 0.302 0.335 0.385 0.408 0.346 0.352 0.345 0.343 0.322 0.400
StableSR [wang2024exploiting] 0.247 0.357 0.364 0.283 0.275 0.295 0.283 0.333 0.294 0.266 0.302 0.340 0.289 0.305 0.316 0.298 0.306 0.234 0.317
Mip-Splatting [Yu2024MipSplatting] 0.275 0.331 0.408 0.288 0.260 0.293 0.296 0.316 0.323 0.250 0.325 0.325 0.281 0.298 0.316 0.284 0.313 0.237 0.345
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.242 0.328 0.367 0.266 0.251 0.277 0.276 0.311 0.292 0.242 0.299 0.321 0.271 0.289 0.304 0.274 0.294 0.221 0.315
Ours + StableSR [wang2024exploiting] 0.248 0.310 0.355 0.246 0.228 0.262 0.259 0.284 0.285 0.225 0.290 0.313 0.257 0.281 0.292 0.253 0.272 0.211 0.300
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
drjohnson playroom bicycle bonsai counter flowers garden kitchen room stump treehill
3DGS (LR) [kerbl3Dgaussians] 0.332 0.339 0.424 0.328 0.306 0.477 0.400 0.312 0.293 0.436 0.482
StableSR [wang2024exploiting] 0.315 0.335 0.407 0.256 0.258 0.461 0.356 0.301 0.267 0.394 0.460
Mip-Splatting [Yu2024MipSplatting] 0.334 0.321 0.318 0.212 0.242 0.397 0.292 0.205 0.245 0.314 0.403
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.311 0.323 0.349 0.232 0.239 0.420 0.315 0.249 0.250 0.368 0.428
Ours + StableSR [wang2024exploiting] 0.300 0.312 0.368 0.238 0.241 0.454 0.328 0.202 0.253 0.371 0.452
Table 9: FID $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
auditorium ignatius palace ballroom panther barn lighthouse playground courtroom m60 temple caterpillar family train francis truck church horse museum
3DGS (LR) [kerbl3Dgaussians] 86.75 85.58 83.25 73.89 40.93 40.48 84.46 82.51 99.19 53.16 81.79 29.93 113.45 43.80 29.64 35.69 58.24 148.48 88.82
StableSR [wang2024exploiting] 68.24 114.03 83.65 49.37 38.50 48.11 72.12 66.81 64.95 40.68 78.22 25.98 68.88 35.61 31.14 32.99 54.69 97.74 54.73
Mip-Splatting [Yu2024MipSplatting] 78.68 52.74 138.29 40.43 24.78 36.49 67.71 59.20 69.94 25.38 82.25 20.50 38.03 30.99 20.36 20.20 48.50 81.54 60.69
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 58.94 79.04 97.00 37.35 26.97 36.74 66.75 54.86 59.58 27.03 75.46 20.10 43.93 29.88 23.64 22.37 45.17 77.02 51.28
Ours + StableSR [wang2024exploiting] 60.48 46.33 70.53 27.26 20.96 27.20 57.93 35.69 54.34 19.79 63.87 16.42 28.28 22.83 16.72 14.82 32.61 58.61 41.93
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
drjohnson playroom bicycle bonsai counter flowers garden kitchen room stump treehill
3DGS (LR) [kerbl3Dgaussians] 60.62 72.04 30.50 72.93 79.04 77.01 47.98 59.89 41.92 89.83 42.08
StableSR [wang2024exploiting] 51.64 67.41 23.55 54.42 64.89 73.04 19.38 49.64 37.09 105.27 40.16
Mip-Splatting [Yu2024MipSplatting] 53.90 50.26 10.21 20.19 36.29 22.76 9.05 12.04 17.57 37.03 20.21
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 47.05 51.10 11.37 26.41 41.87 46.01 10.46 15.43 21.62 42.73 26.52
Ours + StableSR [wang2024exploiting] 43.19 45.09 11.79 22.40 38.03 58.56 10.63 8.47 17.69 50.63 29.83
Table 10: CMMD $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
auditorium ignatius palace ballroom panther barn lighthouse playground courtroom m60 temple caterpillar family train francis truck church horse museum
3DGS (LR) [kerbl3Dgaussians] 1.770 2.234 2.361 1.602 0.718 1.699 1.540 1.925 1.616 1.202 2.522 1.897 3.038 2.153 2.758 2.140 1.705 3.211 2.165
StableSR [wang2024exploiting] 1.093 1.032 1.491 1.073 0.791 1.068 1.092 1.139 0.987 0.907 1.597 1.464 0.647 1.330 1.401 1.143 1.163 0.779 1.142
Mip-Splatting [Yu2024MipSplatting] 1.246 1.046 1.624 0.787 0.633 1.140 0.994 1.151 1.024 0.579 1.853 1.243 0.863 1.384 1.286 1.080 1.029 1.105 1.528
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 1.066 0.931 1.582 0.787 0.516 0.942 0.978 0.998 0.943 0.523 1.795 1.261 0.557 1.383 1.408 1.140 1.049 0.658 1.400
Ours + StableSR [wang2024exploiting] 0.993 1.121 1.568 0.643 0.351 1.134 0.901 1.025 0.951 0.443 1.699 1.330 0.521 1.529 1.355 1.061 0.944 0.897 1.286
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
drjohnson playroom bicycle bonsai counter flowers garden kitchen room stump treehill
3DGS (LR) [kerbl3Dgaussians] 1.003 0.733 0.392 1.011 0.319 1.354 0.242 0.346 0.261 0.926 0.956
StableSR [wang2024exploiting] 0.807 0.714 0.621 0.934 0.440 1.409 0.310 0.441 0.578 0.767 1.023
Mip-Splatting [Yu2024MipSplatting] 0.828 0.552 0.099 0.346 0.195 0.215 0.036 0.191 0.187 0.125 0.257
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.715 0.546 0.138 0.417 0.279 0.782 0.081 0.228 0.378 0.261 0.486
Ours + StableSR [wang2024exploiting] 0.558 0.434 0.159 0.266 0.139 1.108 0.065 0.119 0.184 0.406 0.610
Table 11: DreamSim $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
auditorium ignatius palace ballroom panther barn lighthouse playground courtroom m60 temple caterpillar family train francis truck church horse museum
3DGS (LR) [kerbl3Dgaussians] 0.0969 0.0884 0.1548 0.0881 0.0399 0.0549 0.1154 0.0806 0.1074 0.0552 0.1156 0.0595 0.1215 0.0747 0.0710 0.0620 0.0795 0.1142 0.1207
StableSR [wang2024exploiting] 0.0634 0.0703 0.1497 0.0438 0.0443 0.0613 0.0909 0.0759 0.0566 0.0591 0.0955 0.0516 0.0565 0.0652 0.0555 0.0590 0.0558 0.0505 0.0627
Mip-Splatting [Yu2024MipSplatting] 0.0739 0.0465 0.2459 0.0312 0.0216 0.0392 0.0887 0.0541 0.0592 0.0266 0.1001 0.0392 0.0278 0.0479 0.0341 0.0293 0.0541 0.0438 0.0703
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.0500 0.0511 0.1603 0.0271 0.0264 0.0399 0.0792 0.0564 0.0466 0.0369 0.0936 0.0416 0.0359 0.0504 0.0402 0.0402 0.0459 0.0390 0.0567
Ours + StableSR [wang2024exploiting] 0.0490 0.0381 0.1272 0.0202 0.0187 0.0314 0.0625 0.0369 0.0422 0.0239 0.0720 0.0315 0.0229 0.0410 0.0309 0.0268 0.0331 0.0279 0.0486
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
drjohnson playroom bicycle bonsai counter flowers garden kitchen room stump treehill
3DGS (LR) [kerbl3Dgaussians] 0.0613 0.0534 0.0624 0.0659 0.0453 0.0562 0.0452 0.0454 0.0322 0.0586 0.0558
StableSR [wang2024exploiting] 0.0499 0.0510 0.0348 0.0320 0.0377 0.0598 0.0170 0.0240 0.0345 0.0536 0.0553
Mip-Splatting [Yu2024MipSplatting] 0.0451 0.0345 0.0126 0.0122 0.0166 0.0124 0.0070 0.0046 0.0163 0.0117 0.0252
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.0439 0.0378 0.0150 0.0160 0.0226 0.0331 0.0079 0.0066 0.0212 0.0178 0.0345
Ours + StableSR [wang2024exploiting] 0.0332 0.0327 0.0151 0.0123 0.0177 0.0380 0.0071 0.0037 0.0144 0.0168 0.0364
Table 12: MUSIQ $\uparrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
auditorium ignatius palace ballroom panther barn lighthouse playground courtroom m60 temple caterpillar family train francis truck church horse museum
3DGS (LR) [kerbl3Dgaussians] 55.443 64.965 51.161 56.309 54.610 60.048 48.785 54.510 55.349 53.631 57.565 59.116 65.987 60.573 59.605 61.675 57.991 57.572 62.845
StableSR [wang2024exploiting] 58.573 58.466 49.326 56.501 57.051 60.822 46.028 53.817 55.209 54.985 54.748 60.520 62.883 60.841 55.264 60.637 55.297 54.926 62.311
Mip-Splatting [Yu2024MipSplatting] 51.773 48.745 42.653 50.651 46.800 49.040 33.568 41.815 45.084 47.621 46.236 51.797 49.461 51.300 40.700 47.208 47.433 39.068 53.893
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 57.276 57.690 48.339 54.321 54.969 58.305 44.228 52.796 54.435 53.662 52.762 58.732 61.687 58.721 54.220 57.781 54.378 53.369 61.291
Ours + StableSR [wang2024exploiting] 58.504 60.078 50.762 59.501 56.915 61.660 47.882 58.036 57.351 55.500 55.978 60.757 65.967 60.214 57.281 62.616 57.574 58.244 63.491
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
drjohnson playroom bicycle bonsai counter flowers garden kitchen room stump treehill
3DGS (LR) [kerbl3Dgaussians] 47.916 49.425 54.239 56.108 55.829 58.968 58.961 63.552 50.931 52.085 41.783
StableSR [wang2024exploiting] 52.919 53.324 55.767 59.725 57.965 52.194 54.576 59.952 58.135 44.337 38.276
Mip-Splatting [Yu2024MipSplatting] 38.800 44.105 38.243 44.656 42.445 35.728 44.342 50.600 41.878 35.866 35.112
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 50.849 52.035 54.818 56.237 55.391 53.542 54.318 58.321 55.637 44.039 37.260
Ours + StableSR [wang2024exploiting] 50.594 55.444 57.923 58.501 57.752 51.206 56.236 64.952 55.154 49.551 38.014
Table 13: NIQE $\downarrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]
auditorium ignatius palace ballroom panther barn lighthouse playground courtroom m60 temple caterpillar family train francis truck church horse museum
3DGS (LR) [kerbl3Dgaussians] 4.718 2.628 4.293 2.791 3.884 3.944 4.097 2.744 3.385 3.625 4.081 2.873 2.605 2.994 4.269 2.784 3.169 3.011 2.926
StableSR [wang2024exploiting] 6.531 4.545 5.820 4.335 5.239 5.229 5.415 4.991 4.830 5.022 4.844 4.690 4.214 4.363 5.611 4.361 5.005 4.564 4.356
Mip-Splatting [Yu2024MipSplatting] 5.482 4.693 6.567 4.013 5.614 5.302 5.482 5.086 4.569 4.955 5.039 4.632 5.077 4.449 5.982 5.022 4.653 5.027 4.179
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 6.128 4.187 5.629 3.978 4.904 4.988 5.025 4.537 4.562 4.716 4.573 4.323 4.026 4.137 5.234 4.096 4.644 4.280 4.050
Ours + StableSR [wang2024exploiting] 5.330 3.440 4.720 3.318 3.997 4.657 4.309 3.644 4.029 3.765 4.167 3.340 3.435 3.498 4.559 3.416 3.785 3.632 3.582
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
drjohnson playroom bicycle bonsai counter flowers garden kitchen room stump treehill
3DGS (LR) [kerbl3Dgaussians] 4.760 5.329 2.808 3.964 3.535 2.915 2.741 2.986 4.210 3.137 2.744
StableSR [wang2024exploiting] 5.231 6.225 4.333 5.452 5.533 4.938 5.516 5.633 6.013 4.919 5.105
Mip-Splatting [Yu2024MipSplatting] 6.097 6.426 5.665 6.458 5.730 5.965 6.155 5.600 5.885 5.623 5.462
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 5.208 6.188 4.131 5.228 5.214 4.521 5.074 5.546 5.840 4.846 4.381
Ours + StableSR [wang2024exploiting] 5.054 5.768 3.378 4.309 3.828 3.825 3.949 3.828 4.531 4.052 3.702